A foam cup is filled with a hot water and allowed to cool while being stirred by a paddlewheel.



initially, the water has an internal energy of 200kj,and while cooling it loses 150kj of heat.



the paddlewheel does 25kj of work on water



a) calculate the change in internal energy of water



b) calculate the final internal energy of water




1
Expert's answer
2022-04-05T19:18:49-0400

"\u2206U=150-25=125~kJ,"

"U_2=200-125=75~kJ."